WRITTEN TASK


Quiz 1. 8 Make your own logarithmic function and then construct a graph.


Identify the following:


a.Domain


b.Range


c.X and y – intercepts


d.Zeroes


e.Asymptotes

1
Expert's answer
2021-11-04T18:56:53-0400

Logarithmic function: "y=\log\sqrt{x^2-1}"


a. Domain: "(-\infin,-1)\cup(1,\infin)"


b. Range: "(-\infin,\infin)"


c. x-intercepts: "(-\sqrt{2},0),\;(\sqrt{2},0)"

y-intercept: DNE


d. Zeroes: "x=-\sqrt{2},\;x=\sqrt{2}"


e. Vertical asymptotes: "y=-1,\;y=1"